• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name hoyoung

    Log entry time 00:02:53 on October14,2009

    Entry number 294804

    keyword=Swing shift summary

    Shift workers: Hoyoung Kang(SL), Ge Jin(TO), Maria Leda Sperduto(3rd)

    16:00 Working on Compton measurement was going on from previous shift.
    19:15 Job finished. 10 uA requested
    19:23 spot++ check(run #25043) showed 6mmx5mm raster and correct beam position on empty target
    19:56 spot++ with LH2 target at 10 uA
    19:46 Compton run # 20380 started and production run began at 100 uA. (Compton daq starting was a little bit late.)
    21:47 CODA reboot as ROC 25 was not responding
    22:24 IHWP OUT because we got more than 1M OK pairs (1044k pairs in total). Production restarted

    Run# type curr. rst ok_pairs
    13237 production 100uA 6x5 41k
    13238 production 100uA 6x5 48k
    13240 production 100uA 6x5 26k
    (IHWP OUT)
    13241 production 100uA 6x5 36k
    13242 production 100uA 6x5 (still taking data)